OpenStudy (ohmath123):

Word Problem HELP! Last year, I spent $72 to buy a lot of ping pong balls to use in geometry class. This year, the price of a ping pong ball is 6 cents higher and $72 buys 60 fewer balls. Figure out how many ping- pong balls i bought last year.

OpenStudy (kittiwitti1):

first year: $72.00, x balls second year: $72.00+0.06, x-60 balls

OpenStudy (kittiwitti1):

so let's use the information from the second year to form an equation: ball amount B = money cost C.\[x-60=72.00+0.06\rightarrow x-60=72.06\]Can you do it from here?

OpenStudy (ohmath123):

yes! Thank you @kittiwitti1

OpenStudy (kittiwitti1):

No problem ☺